# A Market for Accuracy: Classification under Competition

### v1

@@ 3 Setup
For a given distribution $p$, if we think of each provider $s_{i}$ as a player and interpret Eq. (3) as their utility, then this defines a game, which we refer to as an accuracy game. The strategy space for each $s_{i}$ is the set of all models in its model class $H_{i}$, and each tuple ${\bm{h}}=(h_{1},\dots,h_{n})$ defines a game state. We will assume the game is played on the empirical distribution induced by ${S}$, but that final payoffs are given by expected market share w.r.t. $p$. Note the game remains well-defined when players have their own ${S}_{i}\sim p_{i}$, although current equilibrium results do not hold for this setting. In terms of information, we work in the full information setting where all players have complete access to the payoff matrix. However, as we will see, optimal strategies for providers require strictly less information.
@@ 4 Analysis
ccc[cell-space-limits=2pt] & $h_{1}$ $h_{2}$
$h_{1}$Block[hvlines]2-2 ${\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\frac{1}{2}a_{1}}},{\leavevmode\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\frac{1}{2}a_{1}}}$ ${\leavevmode\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\frac{1}{2}(a_{1}+\delta_{12})}},{\leavevmode\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\frac{1}{2}(a_{2}+\delta_{21})}}$
@@ 4.1 Warmup: $2\times 2$ accuracy markets
Consider a simple setting with $n=2$ providers and a shared model class of size two, $H=\{h_{1},h_{2}\}$. For example, these could be two available pre-trained models, or an existing model that is already in deployment and a new model that is a possible alternative. Such $2\times 2$ games are fully determined by the tuple $(a_{1},a_{2},\delta_{12},\delta_{21})$—see Table 1. This formulation enables a characterization of all possible equilibria:
@@
In the latter case, the game admits a chicken-like [^4] The standard Chicken game has payoffs of the form Temptation$>$Coordination$>$Neutral$>$Punishment; our variant has T$>$N$>$P$>$C$>0$, which preserves the same PNE. structure: one provider obtains a larger market share by choosing the ‘better’ classifier, while the other must settle for the smaller share. Note that better here does not mean more accurate, as outcomes at equilibrium also depend on discrepancy. The proof of Thm. 1 relies on the following result:
@@ 5 Method
Figure 1: Two-player threshold market. (Left:) Data consists of two class-conditional Gaussians $p(x\,|\,y)={\cal{N}}(ay,\sigma_{y})$ Beginning at $h^{\mathrm{opt}}$, providers compete over who gets the better classifier, $h_{2}$, which secures exclusive access to the larger sector of positive users (blue). (Center:) The game as played over time. Each best response improves market share for both providers, but the second mover ($s_{2}$) prevails. (Right:) Outcomes for increasingly distanced $p(x\,|\,y)$ (here $\sigma_{y}=\sigma$). Equilibrium classifiers are pulled further away, and sacrifice accuracy for increased market share.
@@
where $\ell$ is a proxy loss (e.g., hinge loss or cross-entropy) and ${R}$ is an (optional) regularization term. This approach applies to any type of data and choice of model class $H$. We refer to ${\hat{h}}_{i}$ as the approximate best-response classifier of $s_{i}$. Note that optimizing ${\hat{h}}_{i}$ depends on the other classifiers $h_{-i}$ only through the example weights $w_{i}(x_{j})$. This means that computing the empirical best-response classifier requires only access to the number of other providers that are correct on each data point—not to the actual classifiers in $h_{-i}$. Also, we can observe that the average weight $\frac{1}{|S|}\sum_{i\in S}w_{i}$ has an intuitive meaning: it is the maximum possible market share that could be achieved by a perfect classifier, i.e., when all its predictions are correct. The actual $\mu_{i}$ is then this optimal market share minus the weighted loss incurred by the chosen classifier ${\hat{h}}_{i}$.
@@ 6 Experiments
We now present our empirical investigation of learning in accuracy markets. We begin by demonstrating the basic mechanics of competitive learning on simple synthetic data which allows us to compute best-response classifiers exactly. Then we switch to real data and apply our method from Sec. 5 to accuracy markets across multiple datasets and various learning algorithms. Code is available at https://github.com/Ohadeinav/competition_games.
@@ 6.1 Synthetic data
To gain an understanding of how accuracy markets work, consider a simple setting with $n=2$ providers, binary labels $y\in\{\pm 1\}$, univariate features $x\in\mathbb{R}$ sampled from class-conditional Gaussians $x\sim p(x\,|\,y)={\cal{N}}(ay,\sigma_{y})$, and threshold classifiers $H=\{h_{\tau}(x)=\mathds{1}{\{{x>\tau}\}}\mid\tau\in\mathbb{R}\}$. Figure 1 (left) illustrates this setup for $a=1,\sigma_{-1}=2$, and $\sigma_{+1}=1$, and shows the learned classifiers at equilibrium, $h_{1}$ and $h_{2}$; as Thm. 2 suggests, these are precisely $\rho^{-1}(1/2)$ and $\rho^{-1}(2)$. Note how the region between $h_{1}$ and $h_{2}$ (hatches) is split between the providers: $s_{1}$ is exclusively correct on positive examples, and $s_{2}$ on negatives. The regions to the right of $h_{1}$ and left of $h_{2}$ are shared.
Fig. 1 (center) shows how market shares $\mu_{1},\mu_{2}$ evolve over rounds of best-responses. Here we initialize $h_{1}^{0}=h_{2}^{0}=h^{\mathrm{opt}}$ where $h^{\mathrm{opt}}$ is the optimal classifier (i.e., which maximizes accuracy on $p$). In line with our results from Sec. 4.2, dynamics converge after one round, i.e., each provider responds once, and so $h_{1}=h_{1}^{1}$ and $h_{2}=h_{2}^{1}$. Although $s_{1}$ moves first and improves $\mu_{1}$, this not only improves $\mu_{2}$ for $s_{2}$, but also to a greater extent than that of $s_{1}$. When $s_{2}$ then moves, again both $\mu_{1},\mu_{2}$ increase, but $\mu_{2}$ retains its advantage over $\mu_{1}$. Hence, $s_{2}$ ‘wins’ the chicken game by playing second and obtaining access to the larger exclusive subgroup of positives. Regardless of who wins, users gain from the competition since welfare ($=\mu_{1}+\mu_{2}$) always increases.
Fig. 1 (right) shows how outcomes change for matching gaussians ($\sigma_{+1},\sigma_{-1}=1$) when the class-conditional distributions $p(x\,|\,y=-1)$ and $p(x\,|\,y=1)$ are pulled closer together, achieved by decreasing $a$. When the distributions are far away, $h_{1}$ and $h_{2}$ are at $h^{\mathrm{opt}}$ and so fully share the market. But as overlap increases, several effects take place. First, $h_{1}$ and $h_{2}$ grow further apart and become more distinct in who they target, causing the exclusivity regions to grow in size. Second, since classification becomes harder, the maximal attainable accuracy decreases. The accuracies of $h_{1},h_{2}$ also decrease, but at a faster rate—a result of specialization. Third, we see that welfare—as the sum of market shares— begins at its maximal value of 1 (when the gaussians are perfectly separable), then decreases when the exclusivity doesn’t extend to the tails, and finally returns to 1 when the providers play opposite thresholds. We explore additional aspects on non-matching gaussians in Section D.4.
@@ 6.2 Real data
In terms of market share improvement, we see that competition is helpful for all providers; nonetheless, there can be a large gap between the minimal and maximal improvement. In the COMPAS datasets, this gap begins at smaller values, but grows as $n$ increases. For Adult, whose baseline accuracy is higher, the gap remains mostly stable, but is large to begin with. As competition progresses, the market becomes more concentrated, which also generally increases with $n$. Overall welfare gains are quite high, reaching up to $+65\%$.
@@
Figure 2 shows how the market is partitioned across providers at equilibrium. Here we focus on COMPAS-Arrest with XGBoost and $n=3$ providers; providers are numbered by their order of play (i.e., $s_{1}\prec s_{2}\prec s_{3}$), where this order is preserved across rounds. The plot shows for each provider $s_{i}$ its total market share (bottom left) and accuracy on the entire population (top left) due to its final learned $h_{i}$. The plot also shows the decomposition of the market across all subsets of providers: what proportion is exclusive to $s_{1}$, what is joint to $s_{2}$ and $s_{3}$, what is shared by all, etc. (right). Here we see that $s_{1}$, who moved first, attained the largest market share (36%). However, its accuracy is significantly lower than others, and below 50%. The subsets plot reveals the reason: $s_{1}$ was able to gain exclusive access to 28.5% of the market; it shares an additional $19\%$ with all providers, but only $1\%$ with each of them alone. In contrast, $s_{2}$ and $s_{3}$ share almost all of their users, either as a pair ($44\%$) or along with $s_{1}$. This shows how $s_{1}$ has come to dominate the market by learning a classifier $h_{1}$ that sacrifices accuracy in order to effectively target an exclusive user sector. The low overall accuracy of $h_{1}$ suggests that naïvely optimizing for accuracy without considering the effects of competition can be highly suboptimal in terms of market outcomes.
@@
Whereas our $2\times 2$ analysis from Sec. 4.1 suggested that the game either has a dominant strategy or a chicken-like structure (which implies that moving second is preferable), we see in Fig. 2 that for $n>2$ providers reality is more complex, and in fact moving first allows to dominate the market. Interestingly, this first move induces a 2-player game on the other providers whose equilibrium admits a dominant strategy. To quantify this phenomenon, and assert its robustness across methods, we ran experiments for every combination of datasets and model classes, listed in Appendix C. For each experiment, namely for each combination of model class and dataset, the final market shares were calculated for each provider along with his/her relative position of play, i.e., at what position did the provider perform a best-response. Additionally, each experiment was performed for competitions with 2,3,4,5, and 6 players, so that we can compare dynamics across different market saturations. Figure 3 shows the order of play comparisons, averaged out across all of the experiments that were described above. When the competition game is played with $n=2$ providers, we see a clear preference to be the provider that moves last, characterized by the market share term $\mu_{2}$. The vast majority of experiments showed a significant gain in market share, as seen by the fact that the 25% quantile of values already shows a net-positive gain from moving last. We also note that the expected (mean) competitive advantage of moving last is 3.6% with a median of 4.8%. Given that for 2 players with equal model classes, the market share of a single provider will never exceed $\frac{2}{3}=0.67$ (see Proposition 6), then a $4.8\%$ difference in market share is quite significant.
@@
When the competition game is played with $n\geq 3$ providers, however, we observe an entirely different dynamic. Figure 3 provides two densities: The orange density is as in the 2-provider setting ($\mu_{2}-\mu_{1}$). The blue density is the difference in market share between the player who moved 2nd versus the player who moved 3rd ($\mu_{2}-\mu_{3}$). We find here that the ratios have switched: it is in fact more advantageous to be the provider that moves 1st, as evidenced by the negative orientation of the orange density plot. The next interesting thing that we note is the seeming insignificance of order-of-play beyond the first two positions, as we can observe that the blue density hovers around 0 with relatively low variance. This alludes to the premise put stated above that in competitive settings with 3 or more players, the person who moves 1st is in essence “grabbing his territory”, which then induces all of the other players to play among themselves for the other resources, i.e, consumers.
@@ Appendix A Proofs
$h_{i}=y$ $a_{i}-\delta_{ij}$ $\delta_{ij}$ $=a_{i}$
$h_{i}\neq y$ $\delta_{ji}$ $1-a_{i}-\delta_{ji}$ $=1-a_{i}$
$=a_{j}$ $=1-a_{j}$
@@
Proposition 1 (Market Share).
@@
Proposition 2 (Market share $\leftrightarrow$ accuracy).
@@
From 2, $a_{1}-\delta_{12}=a_{2}-\delta_{21}$, meaning: $a_{1}=a_{2}+\delta_{12}-\delta_{21}$. Therefore:
@@
We will now prove Proposition 2:
@@
From Proposition 1, We know: $\mu(h_{1}|h_{2})=a_{1}+\delta_{12}$, and $\mu(h_{2}|h_{1})=a_{2}+\delta_{21}$.
@@
Where the last inequality follows from 1. ∎
Lemma 1.
@@
From Proposition 1 we know that $U(h_{1}|h_{1})=\frac{1}{2}a_{1},U(h_{2}|h_{2})=\frac{1}{2}a_{2},U(h_{i}|h_{j})=\frac{1}{2}(a_{i}+\delta_{ij})$. Therefore,
@@
Where the substitution of the RHS in the 2nd line comes from 2.
@@
Using 2, the proof of Lemma 1 is almost immediate: from 2 we know that $a_{1}-a_{2}=\delta_{12}-\delta_{21}$.
The 2 players will choose differing strategies in the 2x2 game if and only if $U(h_{1}|h_{2})>U(h_{2}|h_{2})\ \textbf{and}\ U(h_{2}|h_{1})>U(h_{1}|h_{1})$, which holds if and only if $\delta_{h_{1}}>\frac{1}{2}\delta_{h_{2}}\textbf{and}\ \delta_{h_{2}}>\frac{1}{2}\delta_{h_{1}}$, which holds if and only if $|\delta_{h_{1}}-\delta_{h_{2}}|<\frac{1}{3}(\delta_{12}+\delta_{21})$. Since we know $a_{1}-a_{2}=\delta_{12}-\delta_{21}$, this proves Lemma 1.
Theorem 1 (2x2 PNEs.)
From the inequalities in Lemma 1, we receive the conditions for which the providers play anti-coordinated strategies.
@@
Theorem 2 (Threshold best-responses under MLR).
@@
[…]
@@
Immediate from the fact that the strategy space of both players can be reduced to the 2 candidate best-responses stipulated in Theorem 2. ∎
Theorem 3 (General threshold best-responses).
@@
As in the proof of Theorem 2, We will split $[a,b]$ into sub-intervals $[a,h],[h,b]$ and calculate the set of possible best responses for each interval:
@@
As explained in the proof of Theorem 2, to calculate the best response in $[a,h]$,it is sufficient to maximize $U_{[a,h]}$; and to calculate the best-response in $[h,b]$, it is sufficient to maximize $U_{[h,b]}$.
@@
Since we are given that in the segment $[h_{1},x],\ g>\frac{1}{2}$, then it folows that $\forall u,f_{1}(u)-\frac{1}{2}f_{0}(u)>0$, and therefore the integral must be positive and hence $U_{[a,h]}(h_{1}|h)-U_{[a,h]}(x|h)>0$.
@@
We will show that $U_{[a,h]}(h_{2}|h)-U_{[a,h]}(x|h)>0$, in a similar manner to 3:
@@
Since we are given that in the segment $[x,h_{2}],\ g<\frac{1}{2}$, then it folows that $\forall u,f_{0}(u)-\frac{1}{2}f_{1}(u)>0$, and therefore the integral must be positive and hence $U_{[a,h]}(h_{2}|h)-U_{[a,h]}(x|h)>0$.
@@
Let $(x,y)\subseteq[a,h]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})>\frac{1}{2}$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(x)=\frac{1}{2}$ and $g^{\prime}(x)>0$, or $x=a$. From 3, we receive that $U_{[a,h]}(x|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.
@@
Let $(x,y)\subseteq[a,h]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})<\frac{1}{2}$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(y)=\frac{1}{2}$ and $g^{\prime}(x)>0$, or $y=h$. From 4, we receive that $U_{[a,h]}(y|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.
@@
Since we are given that in the segment $[x,h_{2}],\ g>2$, then it follows that $\forall u,\frac{1}{2}f_{1}(u)-f_{0}(u)>0$, and therefore the integral must be positive and hence $U_{[h,b]}(h_{1}|h)-U_{[h,b]}(x|h)>0$. ∎
@@
Since we are given that in the segment $[x,h_{2}],\ g<2$, then it follows that $\forall u,f_{0}(u)-\frac{1}{2}f_{1}(u)>0$, and therefore the integral must be positive and hence $U_{[h,b]}(h_{2}|h)-U_{[h,b]}(x|h)>0$.
@@
Let $(x,y)\subseteq[a,b]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})>2$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(x)=2$ and $g^{\prime}(x)>0$, or $x=a$. From 5, we receive that $U_{[a,h]}(x|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.

### v2

@@ 3 Setup
For a given distribution $p$, if we think of each provider $s_{i}$ as a player and interpret Eq. (3) as their utility, then this defines a game, which we refer to as an accuracy game. The strategy space for each $s_{i}$ is the set of all models in its model class $H_{i}$, and each tuple ${\bm{h}}=(h_{1},\dots,h_{n})$ defines a game state. We will assume the game is played on the empirical distribution induced by S{S}, but that final payoffs are given by expected market share w.r.t. $p$. Note the game remains well-defined when players have their own ${S}_{i}\sim p_{i}$, although current equilibrium results do not hold for this setting. In terms of information, we work in the full information setting where all players have complete access to the payoff matrix. However, as we will see, optimal strategies for providers require strictly less information.
@@ 4 Analysis
ccc[cell-space-limits=2pt] $h_{1}$ $h_{2}$
[…]
@@ 4.1 Warmup: $2\times 2$ accuracy markets
Consider a simple setting with $n=2$ providers and a shared model class of size two, $H=\{h_{1},h_{2}\}$. For example, these could be two available pre-trained models, or an existing model that is already in deployment and a new model that is a possible alternative. Such $2\times 2$ games are fully determined by the tuple $(a_{1},a_{2},\delta_{12},\delta_{21})$—see Table˜1. This formulation enables a characterization of all possible equilibria:
@@
In the latter case, the game admits a chicken-like [^4] The standard Chicken game has payoffs of the form Temptation>>Coordination>>Neutral>>Punishment; our variant has T>>N>>P>>C>0>0, which preserves the same PNE. structure: one provider obtains a larger market share by choosing the ‘better’ classifier, while the other must settle for the smaller share. Note that better here does not mean more accurate, as outcomes at equilibrium also depend on discrepancy. The proof of Thm. 1 relies on the following result:
@@ 5 Method
Figure 1: Two-player threshold market. (Left:) Data consists of two class-conditional Gaussians $p(x\,|\,y)={\cal{N}}(ay,\sigma_{y})$ Beginning at $h^{\mathrm{opt}}$, providers compete over who gets the better classifier, $h_{2}$, which secures exclusive access to the larger sector of positive users (blue). (Center:) The game as played over time. Each best response improves market share for both providers, but the second mover ($s_{2}$) prevails. (Right:) Outcomes for increasingly distanced p(x|y)p(x\,|\,y) (here $\sigma_{y}=\sigma$). Equilibrium classifiers are pulled further away, and sacrifice accuracy for increased market share.
@@
where $\ell$ is a proxy loss (e.g., hinge loss or cross-entropy) and R{R} is an (optional) regularization term. This approach applies to any type of data and choice of model class $H$. We refer to ${\hat{h}}_{i}$ as the approximate best-response classifier of $s_{i}$. Note that optimizing ${\hat{h}}_{i}$ depends on the other classifiers $h_{-i}$ only through the example weights $w_{i}(x_{j})$. This means that computing the empirical best-response classifier requires only access to the number of other providers that are correct on each data point—not to the actual classifiers in $h_{-i}$. Also, we can observe that the average weight $\frac{1}{|S|}\sum_{i\in S}w_{i}$ has an intuitive meaning: it is the maximum possible market share that could be achieved by a perfect classifier, i.e., when all its predictions are correct. The actual $\mu_{i}$ is then this optimal market share minus the weighted loss incurred by the chosen classifier ${\hat{h}}_{i}$.
@@ 6 Experiments
We now present our empirical investigation of learning in accuracy markets. We begin by demonstrating the basic mechanics of competitive learning on simple synthetic data which allows us to compute best-response classifiers exactly. Then we switch to real data and apply our method from Sec. 5 to accuracy markets across multiple datasets and various learning algorithms. Code is publicly available at https://github.com/BML-Technion/market4acc.
@@ 6.1 Synthetic data
To gain an understanding of how accuracy markets work, consider a simple setting with $n=2$ providers, binary labels $y\in\{\pm 1\}$, univariate features $x\in\mathbb{R}$ sampled from class-conditional Gaussians $x\sim p(x\,|\,y)={\cal{N}}(ay,\sigma_{y})$, and threshold classifiers $H=\{h_{\tau}(x)=\mathds{1}{\{{x>\tau}\}}\mid\tau\in\mathbb{R}\}$. Figure˜1 (left) illustrates this setup for $a=1,\sigma_{-1}=2$, and $\sigma_{+1}=1$, and shows the learned classifiers at equilibrium, $h_{1}$ and $h_{2}$; as Thm. 2 suggests, these are precisely $\rho^{-1}(1/2)$ and $\rho^{-1}(2)$. Note how the region between $h_{1}$ and $h_{2}$ (hatches) is split between the providers: $s_{1}$ is exclusively correct on positive examples, and $s_{2}$ on negatives. The regions to the right of $h_{1}$ and left of $h_{2}$ are shared.
Fig. 1 (center) shows how market shares $\mu_{1},\mu_{2}$ evolve over rounds of best-responses. Here we initialize $h_{1}^{0}=h_{2}^{0}=h^{\mathrm{opt}}$ where $h^{\mathrm{opt}}$ is the optimal classifier (i.e., which maximizes accuracy on $p$). In line with our results from Sec. 4.2, dynamics converge after one round, i.e., each provider responds once, and so $h_{1}=h_{1}^{1}$ and $h_{2}=h_{2}^{1}$. Although $s_{1}$ moves first and improves $\mu_{1}$, this not only improves $\mu_{2}$ for $s_{2}$, but also to a greater extent than that of $s_{1}$. When $s_{2}$ then moves, again both $\mu_{1},\mu_{2}$ increase, but $\mu_{2}$ retains its advantage over $\mu_{1}$. Hence, $s_{2}$ ‘wins’ the chicken game by playing second and obtaining access to the larger exclusive subgroup of positives. Regardless of who wins, users gain from the competition since welfare ($\mu_{1}+\mu_{2}$) always increases.
Fig. 1 (right) shows how outcomes change for matching gaussians ($\sigma_{+1},\sigma_{-1}=1$) when the class-conditional distributions p(x|y=1)p(x\,|\,y=-1) and p(x|y=1)p(x\,|\,y=1) are pulled closer together, achieved by decreasing $a$. When the distributions are far away, $h_{1}$ and $h_{2}$ are at $h^{\mathrm{opt}}$ and so fully share the market. But as overlap increases, several effects take place. First, $h_{1}$ and $h_{2}$ grow further apart and become more distinct in who they target, causing the exclusivity regions to grow in size. Second, since classification becomes harder, the maximal attainable accuracy decreases. The accuracies of $h_{1},h_{2}$ also decrease, but at a faster rate—a result of specialization. Third, we see that welfare—as the sum of market shares— begins at its maximal value of 1 (when the gaussians are perfectly separable), then decreases when the exclusivity doesn’t extend to the tails, and finally returns to 1 when the providers play opposite thresholds. We explore additional aspects on non-matching gaussians in Section˜D.4.
@@ 6.2 Real data
In terms of market share improvement, we see that competition is helpful for all providers; nonetheless, there can be a large gap between the minimal and maximal improvement. In the COMPAS datasets, this gap begins at smaller values, but grows as $n$ increases. For Adult, whose baseline accuracy is higher, the gap remains mostly stable, but is large to begin with. As competition progresses, the market becomes more concentrated, which also generally increases with $n$. Overall welfare gains are quite high, reaching up to $65\%$.
@@
Figure˜2 shows how the market is partitioned across providers at equilibrium. Here we focus on COMPAS-Arrest with XGBoost and $n=3$ providers; providers are numbered by their order of play (i.e., $s_{1}\prec s_{2}\prec s_{3}$), where this order is preserved across rounds. The plot shows for each provider $s_{i}$ its total market share (bottom left) and accuracy on the entire population (top left) due to its final learned $h_{i}$. The plot also shows the decomposition of the market across all subsets of providers: what proportion is exclusive to $s_{1}$, what is joint to $s_{2}$ and $s_{3}$, what is shared by all, etc. (right). Here we see that $s_{1}$, who moved first, attained the largest market share (36%). However, its accuracy is significantly lower than others, and below 50%. The subsets plot reveals the reason: $s_{1}$ was able to gain exclusive access to 28.5% of the market; it shares an additional $19\%$ with all providers, but only $1\%$ with each of them alone. In contrast, $s_{2}$ and $s_{3}$ share almost all of their users, either as a pair ($44\%$) or along with $s_{1}$. This shows how $s_{1}$ has come to dominate the market by learning a classifier $h_{1}$ that sacrifices accuracy in order to effectively target an exclusive user sector. The low overall accuracy of $h_{1}$ suggests that naïvely optimizing for accuracy without considering the effects of competition can be highly suboptimal in terms of market outcomes.
@@
Whereas our $2\times 2$ analysis from Sec. 4.1 suggested that the game either has a dominant strategy or a chicken-like structure (which implies that moving second is preferable), we see in Fig. 2 that for $n>2$ providers reality is more complex, and in fact moving first allows to dominate the market. Interestingly, this first move induces a 2-player game on the other providers whose equilibrium admits a dominant strategy. To quantify this phenomenon, and assert its robustness across methods, we ran experiments for every combination of datasets and model classes, listed in Appendix˜C. For each experiment, namely for each combination of model class and dataset, the final market shares were calculated for each provider along with his/her relative position of play, i.e., at what position did the provider perform a best-response. Additionally, each experiment was performed for competitions with 2,3,4,5, and 6 players, so that we can compare dynamics across different market saturations. Figure˜3 shows the order of play comparisons, averaged out across all of the experiments that were described above. When the competition game is played with $n=2$ providers, we see a clear preference to be the provider that moves last, characterized by the market share term $\mu_{2}$. The vast majority of experiments showed a significant gain in market share, as seen by the fact that the 25% quantile of values already shows a net-positive gain from moving last. We also note that the expected (mean) competitive advantage of moving last is 3.6% with a median of 4.8%. Given that for 2 players with equal model classes, the market share of a single provider will never exceed $\frac{2}{3}=0.67$ (see Proposition˜6), then a $4.8\%$ difference in market share is quite significant.
@@
When the competition game is played with $n\geq 3$ providers, however, we observe an entirely different dynamic. Figure˜3 provides two densities: The orange density is as in the 2-provider setting ($\mu_{2}-\mu_{1}$). The blue density is the difference in market share between the player who moved 2nd versus the player who moved 3rd ($\mu_{2}-\mu_{3}$). We find here that the ratios have switched: it is in fact more advantageous to be the provider that moves 1st, as evidenced by the negative orientation of the orange density plot. The next interesting thing that we note is the seeming insignificance of order-of-play beyond the first two positions, as we can observe that the blue density hovers around 0 with relatively low variance. This alludes to the premise put stated above that in competitive settings with 3 or more players, the person who moves 1st is in essence “grabbing his territory”, which then induces all of the other players to play among themselves for the other resources, i.e, consumers.
@@ Appendix A Proofs
$h_{i}=y$ $a_{i}-\delta_{ij}$ $\delta_{ij}$ $a_{i}$
$h_{i}\neq y$ $\delta_{ji}$ $1-a_{i}-\delta_{ji}$ $1-a_{i}$
$a_{j} =$$1-a_{j}$
@@
Proposition˜1 (Market Share).
@@
Proposition˜2 (Market share $\leftrightarrow$ accuracy).
@@
From ˜2, $a_{1}-\delta_{12}=a_{2}-\delta_{21}$, meaning: $a_{1}=a_{2}+\delta_{12}-\delta_{21}$. Therefore:
@@
We will now prove Proposition˜2:
@@
From Proposition˜1, We know: $\mu(h_{1}|h_{2})=a_{1}+\delta_{12}$, and $\mu(h_{2}|h_{1})=a_{2}+\delta_{21}$.
@@
Where the last inequality follows from ˜1. ∎
Lemma˜1.
@@
From Proposition˜1 we know that $U(h_{1}|h_{1})=\frac{1}{2}a_{1},U(h_{2}|h_{2})=\frac{1}{2}a_{2},U(h_{i}|h_{j})=\frac{1}{2}(a_{i}+\delta_{ij})$. Therefore,
@@
Where the substitution of the RHS in the 2nd line comes from ˜2.
@@
Using ˜2, the proof of Lemma˜1 is almost immediate: from ˜2 we know that $a_{1}-a_{2}=\delta_{12}-\delta_{21}$.
The 2 players will choose differing strategies in the 2x2 game if and only if $U(h_{1}|h_{2})>U(h_{2}|h_{2})\ \textbf{and}\ U(h_{2}|h_{1})>U(h_{1}|h_{1})$, which holds if and only if $\delta_{h_{1}}>\frac{1}{2}\delta_{h_{2}}\textbf{and}\ \delta_{h_{2}}>\frac{1}{2}\delta_{h_{1}}$, which holds if and only if $|\delta_{h_{1}}-\delta_{h_{2}}|<\frac{1}{3}(\delta_{12}+\delta_{21})$. Since we know $a_{1}-a_{2}=\delta_{12}-\delta_{21}$, this proves Lemma˜1.
Theorem˜1 (2x2 PNEs.)
From the inequalities in Lemma˜1, we receive the conditions for which the providers play anti-coordinated strategies.
@@
Theorem˜2 (Threshold best-responses under MLR).
@@
This is a straightforward expression of the utility. Any points outside $h_{1},h]$ have an identical classification for both $h$ and $h_{1}$, and so the only difference in utility is found inside the interval $h_{1},h]$, which $h_{1}$ classifies as positive and $h$ classifies as negative. Therefore the gain in utility is $\int_{h_{1}}^{h}f_{1}(x)dx$, but the loss on the negative points is $\frac{1}{2}\int_{h_{1}}^{h}f_{0}(x)dx$, since the utility on those points would be otherwise shared with $h$.
@@
Immediate from the fact that the strategy space of both players can be reduced to the 2 candidate best-responses stipulated in Theorem˜2. ∎
Theorem˜3 (General threshold best-responses).
@@
As in the proof of Theorem˜2, We will split $[a,b]$ into sub-intervals $[a,h],[h,b]$ and calculate the set of possible best responses for each interval:
@@
As explained in the proof of Theorem˜2, to calculate the best response in $[a,h]$,it is sufficient to maximize $U_{[a,h]}$; and to calculate the best-response in $[h,b]$, it is sufficient to maximize $U_{[h,b]}$.
@@
Since we are given that in the segment $h_{1},x],\ g>\frac{1}{2}$, then it folows that $\forall u,f_{1}(u)-\frac{1}{2}f_{0}(u)>0$, and therefore the integral must be positive and hence $U_{[a,h]}(h_{1}|h)-U_{[a,h]}(x|h)>0$.
@@
We will show that $U_{[a,h]}(h_{2}|h)-U_{[a,h]}(x|h)>0$, in a similar manner to ˜3:
@@
Since we are given that in the segment $x,h_{2}],\ g<\frac{1}{2}$, then it folows that $\forall u,f_{0}(u)-\frac{1}{2}f_{1}(u)>0$, and therefore the integral must be positive and hence $U_{[a,h]}(h_{2}|h)-U_{[a,h]}(x|h)>0$.
@@
Let $(x,y)\subseteq[a,h]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})>\frac{1}{2}$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(x)=\frac{1}{2}$ and $g^{\prime}(x)>0$, or $x=a$. From ˜3, we receive that $U_{[a,h]}(x|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.
@@
Let $(x,y)\subseteq[a,h]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})<\frac{1}{2}$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(y)=\frac{1}{2}$ and $g^{\prime}(x)>0$, or $y=h$. From ˜4, we receive that $U_{[a,h]}(y|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.
@@
Since we are given that in the segment $x,h_{2}],\ g>2$, then it follows that $\forall u,\frac{1}{2}f_{1}(u)-f_{0}(u)>0$, and therefore the integral must be positive and hence $U_{[h,b]}(h_{1}|h)-U_{[h,b]}(x|h)>0$. ∎
@@
Since we are given that in the segment $x,h_{2}],\ g<2$, then it follows that $\forall u,f_{0}(u)-\frac{1}{2}f_{1}(u)>0$, and therefore the integral must be positive and hence $U_{[h,b]}(h_{2}|h)-U_{[h,b]}(x|h)>0$.
@@
Let $(x,y)\subseteq[a,b]$ be the largest consecutive interval that includes $h_{1}$ such that $\forall h^{\prime}\in(x,y)\to g(h^{\prime})>2$. Since $f_{0},f_{1}$ are continuous, then we know $g$ is continuous. Therefore, either $g(x)=2$ and $g^{\prime}(x)>0$, or $x=a$. From ˜5, we receive that $U_{[a,h]}(x|h)>U_{[a,h]}(h_{1}|h)$, and therefore $h_{1}$ cannot be a best response.